suppose the federal reserve injects 100 million into financial system with a 25% required reserves ratio. what is the maximum total increase in the money supply from this policy?

A- $100 m

B-$250 m

C-$400 m

D- $500m
Answer #1

Increase in money supply=(1/reserve ratio)*the amount fed injects in the economy.

I/RR=1/25%

Money multiplier=1/RR=4

increase in money supply=4*100=400

Answer-C